Suppose you travel north for 35 kilometers then travel east 65 kilometers. How far are you from your starting point? North and east can be considered the directions of the y-and x-axis respectively. Round to the tenth's place.
Answer by checkley71(8403)

35^2+65^2=D^2
1225+4225=D^2
5450=D^2
D=73.824115 KM. FROM THE STARTING POINT.